### v1

# Hybrid Quantum Recurrent Neural Network For Remaining Useful Life Prediction

Abstract
Predictive maintenance in aerospace heavily relies on accurate estimation of the remaining useful life of jet engines. In this paper, we introduce a Hybrid Quantum Recurrent Neural Network framework, combining Quantum Long Short-Term Memory layers with classical dense layers for Remaining Useful Life forecasting on NASA’s Commercial Modular Aero-Propulsion System Simulation dataset. Each Quantum Long Short-Term Memory gate replaces conventional linear transformations with Quantum Depth-Infused circuits, allowing the network to learn high-frequency components more effectively. Experimental results demonstrate that, despite having fewer trainable parameters, the Hybrid Quantum Recurrent Neural Network achieves up to a 5% improvement over a Recurrent Neural Network based on stacked Long Short-Term Memory layers in terms of mean root mean squared error and mean absolute error. Moreover, a thorough comparison of our method with established techniques, including Random Forest, Convolutional Neural Network, and Multilayer Perceptron, demonstrates that our approach, which achieves a Root Mean Squared Error of 15.46, surpasses these baselines by approximately 13.68%, 16.21%, and 7.87%, respectively. Nevertheless, it remains outperformed by certain advanced joint architectures. Our findings highlight the potential of hybrid quantum-classical approaches for robust time-series forecasting under limited data conditions, offering new avenues for enhancing reliability in predictive maintenance tasks.
Remaining Useful Life, Quantum Machine Learning, Recurrent Neural Network, LSTM, Predictive Maintenance, Time-Series Forecasting
preprint: APS/123-QED
I Introduction
Accurate estimation of the remaining useful life (RUL) of critical machinery is a cornerstone of modern reliability and risk-management strategies si2011remaining ; lee2014prognostics ; berghout2022systematic . Nowhere is this more evident than in the aviation industry: the timely prediction of gas turbine jet engine failures not only bolsters safety by preventing catastrophic breakdowns zhang2016multiobjective , but also serves to streamline maintenance and resource allocation kang2021remaining . Indeed, robust forecasting models enable operators to defer costly repairs until necessary while avoiding the risks of overdue overhauls, ultimately leading to reduced downtime and more efficient fleet utilization huang2019bidirectional ; ferreira2022remaining .
Time-series forecasting methods for RUL estimation can be broadly categorized into statistical approaches and machine learning techniques. Traditional statistical methods, including autoregressive, moving average, and autoregressive integrated moving average models, typically excel under assumptions of stationarity and linearity box2015time ; shumway2000time . These models are prized for their interpretability and rigorous theoretical foundation but can struggle to capture the nonlinear and interdependent patterns that characterize real-world industrial and aerospace data zhang2003time .
Machine learning (ML) has exhibited remarkable success in tackling complex predictive tasks across numerous fields, including economics stock2006forecasting , healthcare ML2023Healthcare , energy ML2025Energy , meteorology zhou2019forecasting , and industry Ml2025Industry . Within this landscape, recurrent neural networks (RNNs) stand out due to their ability to capture temporal dependencies in data. However, they are often hindered by vanishing and exploding gradients bengio1994learning , which complicate model training. Long short-term memory (LSTM) networks hochreiter1997long ; gers2000learning ; graves2013speech ; sak2014long alleviate these issues by effectively modeling nonlinear interactions, long-term trends, and seasonal or cyclical patterns, thus finding broad applicability in time series tasks, including RUL estimation LSTM2024Summary .
Despite these achievements, classical ML strategies can be severely weakened by limited or noisy data, high-dimensional feature spaces, and intricate fault dynamics Bishop2007Pattern ; Goodfellow2016Deep . In recent years, quantum computing has progressed from theoretical constructs to initial practical demonstrations, leveraging quantum properties such as entanglement and superposition to enable computational processes beyond the reach of traditional hardware Nielsen2011Quantum ; biamonte2017quantum . This development has led to quantum algorithms designed for optimization, simulation, and cryptography, thereby paving the way for more specialized implementations montanaro2016quantum ; Preskill2018quantumcomputingin .
Finally, the new tool that allows the use of the advantages of ML algorithms and quantum computations emerged — quantum machine learning (QML), which proved its usability by solving problems typified by constrained datasets or high-dimensional feature spaces rebentrost2014quantum ; ciliberto2018quantum ; schuld2018supervised ; cao2017quantum ; Alharbi2024QCNN . QML utilises high-dimensional Hilbert spaces to encode input features, allowing it to detect more complex patterns with fewer parameters than its classical counterpart havlivcek2019supervised ; schuld2019quantum . This proved to be especially useful when applied to problems with nonstationary or noisy signals, which are frequently encountered in industrial maintenance and prognostic applications  sagingalieva2023photovoltaic ; emmanoulopoulos2022quantummachinelearningfinance ; protasevich2025multi ; lee2025predictive .
A promising approach for leveraging these quantum advantages involves hybrid quantum-classical neural networks (HQNNs), which combine classical and quantum ML in one single framework arthur2022hybridquantumclassicalneuralnetwork ; sagingalieva2023photovoltaic ; sagingalieva2023hybrid ; Bischof2025Hybrid ; SUN2025130226 . Inside this structure, the quantum component plays a specific role in data processing or feature encoding while leaving computationally hard tasks like backpropagation and parameter optimization for the classical routines. By selectively introducing quantum effects, HQNNs retain the scalability and reliability of conventional deep learning pipelines broughton2021tensorflowquantumsoftwareframework . Preliminary studies indicate that HQNNs can match or exceed the performance of their classical counterparts, often exhibiting greater resilience against overfitting abbas2021power ; Berberich2024training . The inherent parameter efficiency of HQNNs further enhances their suitability for scenarios involving complex, high-dimensional inputs or limited data availability — situations frequently observed in predictive maintenance sagingalieva2023hybrid ; anoshin2024hybrid .
In this paper, we propose a novel Hybrid Quantum Recurrent Neural Network (HQRNN) tailored for jet engine RUL forecasting. Our approach replaces the linear transformations in each gate of a classical LSTM with a Quantum Depth-Infused (QDI) circuit sagingalieva2023hybrid ; anoshin2024hybrid ; lusnig2024hybrid , creating a quantum-enhanced LSTM (QLSTM) chen2022quantum ; sagingalieva2023photovoltaic . By encoding sensor measurements in a quantum Hilbert space, these QDI layers effectively capture complex temporal patterns while requiring fewer trainable parameters than a purely classical baseline chen2022quantum . We evaluate our method on the NASA Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset saxena2008turbofan , a widely used benchmark in turbofan engine degradation studies, and demonstrate improved predictive performance in terms of both root mean squared error (RMSE) and mean absolute error (MAE). Moreover, experimental results suggest that our hybrid approach can help mitigate the challenges associated with limited data, underscoring its potential benefits in industrial time-series applications sagingalieva2023hybrid .
The C-MAPSS dataset is a widely used benchmark for predicting RUL of aircraft engines. Recent studies addressing this prediction problem include approaches based on Temporal Convolutional Neural Networks (TCNN) combined with Transformers wang2021remaining and joint CNN-LSTM models enhanced with attention mechanisms deng2024prediction ; both articles also provide comprehensive summaries of alternative methods. In Section VI, we present a detailed comparative analysis of our proposed HQRNN against classical machine learning techniques and simple neural network architectures (Table 2), as well as recent state-of-the-art joint deep learning approaches (Table 3). Our HQRNN approach demonstrates superior performance compared to traditional ML and simple neural network models, though it currently underperforms relative to several advanced joint models. These results highlight promising opportunities for future research focused on integrating quantum-enhanced recurrent structures with complementary deep learning methods.
The rest of the paper is structured as follows: Section II describes the C-MAPSS dataset used in this study. Section III outlines our proposed HQRNN architecture and the underlying QDI circuits. We detail the experimental setup and report empirical findings in Section IV, followed by a deeper analysis of the quantum circuit properties in Section V. Lastly, Section VI concludes with a discussion of the implications and future directions for quantum-enhanced RUL forecasting.
II Dataset
The NASA C-MAPSS dataset is a widely used multivariate time series for analyzing gas turbine engine degradation. In this work, we employ the “FD001” subset, which includes run-to-failure data for a fleet of 100 engines. Each engine progresses from a nominal state to the point of failure, providing sensor measurements over multiple cycles. The dataset comprises four main feature groups:
  1. 1.
    An engine identifier (ranging from 1 to 100),
  2. 2.
    A time index in cycles,
  3. 3.
    Three operational settings,
  4. 4.
    Twenty-one sensor measurements.
For simplicity, we exclude the operational settings, focusing instead on the sensor measurements for predicting RUL.
The data are split into a training set and a test set. In the training set, each engine’s sensor measurements are recorded until failure occurs, whereas in the test set, the measurements end at an arbitrary cycle prior to failure. So the goal is to predict the number of remaining operational cycles before failure in the test set, i.e., the number of operational cycles after the last cycle that the engine will continue to operate. A vector of true RUL values for the test data is provided.
We design our predictive model as follows. A model gets a fixed size window of consecutive engine cycles as input and yields a single estimated RUL value as output. This means that for each entry in the training dataset there should be a corresponding target RUL value.
Creation of these RUL values depends on the degradation model. A linear degradation model, for instance, linearly decreases the RUL value to zero as engine approaches the point of failure. In this work, however, we will use a piecewise linear degradation model, which follows the idea that for some time an engine is operating normally (RUL is constant, also known as “early RUL”) and at some point develops a fault, which leads to linear decrease of its useful life, see Figure 1. We chose to set early RUL to 125 cycles, based on the minimum run length of 128 cycles observed in the training data.
Figure 1: Remaining useful life of engine 2 according to linear degradation and piecewise linear degradation models. This work adopts the piecewise linear model with an early RUL threshold of 125.
Upon examining the training data, we found that some sensor features remain constant over all cycles; hence, we exclude these as they contribute no predictive information. After removing redundant features, 14 sensor channels remain. We then apply a standard scaling transformation to these channels, using only the training data to fit the scaler.
Because our proposed architecture relies on LSTM-like layers for RUL prediction, we segment each time series into windows of length 30 cycles. Each window is therefore a 30 $\times$ 14 matrix of sensor values. The network takes this window as input and produces an RUL estimate for the cycle immediately following the window, facilitating a supervised learning setup for the engine-degradation problem.
III Hybrid Quantum Recurrent Neural Network
In this work, we employ a HQRNN to predict RUL of jet engines. The core of the HQRNN architecture is a stack of QLSTM layers, followed by a series of fully connected layers for final regression. QLSTM is derived from the conventional LSTM by substituting the linear transformations in each of the four gates (forget, input, update, and output) with a QDI layer.
Classical deep neural networks often bias their learning toward lower-frequency components, a phenomenon sometimes referred to as the “F-Principle” xu2019training . By contrast, the QDI circuits used within QLSTM inherently operate in a Fourier-transformed space, allowing them to capture higher-frequency components more effectively schuld2021effect . This property is particularly advantageous for time-series data, as fine-grained temporal fluctuations can be challenging for purely classical models to capture.
Figure 2: (a) HQRNN model pipeline. A data window of size $W\times 14$ (where $W$ is the window size and 14 is the number of sensor features) is processed through three stacked QLSTM layers followed by classical Dense layers, yielding a single RUL value. The dimensions of the QLSTM layers are 32, 16, and 8, while the Dense layers transition from $8\times W$ to 16, 16 to 32, and finally 32 to 1. (b) Structure of the QLSTM layer. Conventional linear transformations are replaced by QDI layers for each of the four LSTM gates (forget, input, update, output). (c) A schematic of the QDI layer used in the QLSTM. Each input feature is encoded via parameterized $R_{z}$ gates on a 4-qubit quantum circuit. The variation part consists of parameterized rotations of $R_{z}$ and CNOT gates. The blue block repeats $n=1$ time. The observable is the Pauli $Y$ matrix.
Figure 2 provides an overview of the HQRNN model and its constituent layers. Fig. 2(a) shows the complete pipeline: an input window of sensor measurements of size $W\times 14$ is passed through three stacked QLSTM layers with latent dimensions of 32, 16, and 8, respectively. The output of the QLSTM stack is then fed into classical Dense layers, which reduce the dimension from $8\times W$ to 16, then 32, and finally to a single value representing the predicted RUL. Fig. 2(b) depicts the internal structure of a single QLSTM layer, highlighting the replacement of each conventional linear gate transformation with a individual QDI layer.
Fig. 2(c) shows the detailed design of the QDI layer itself. Before encoding the input data, a parameterized $R_{x}$ rotation is applied to the initial state of the four qubits using the trained parameters as rotation angles, these are combined with a ring of controlled NOT (CNOT) gates (orange operation block).
[…]
$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix}$
This is followed by an encoding block represented by $R_{z}$ rotations using input features as rotation angles (blue rectangles).
$R_{z}(\phi)=\begin{pmatrix}\text{exp}(-i\frac{\phi}{2})&0\\ 0&\text{exp}(i\frac{\phi}{2})\end{pmatrix}$
This is followed again by the variation part $R_{x}$ combined with CNOT gates. The blue operation block is repeated $n=1$ time. The quantum circuit is completed by measurement in the basis of eigenvectors of the Pauli $Y$ matrix.
For comparison, we also train a purely classical LSTM-based RNN that retains the same overall structure but replaces the QDI layers with standard linear transformations.
IV Training and Results
In our experiments, the proposed HQRNN model takes a $30\times 14$ window of sensor data and outputs the corresponding RUL. We partition the training dataset such that $20\%$ of it is used for validation, while the test dataset is used exclusively for final performance assessment. The training objective is to minimize the mean squared error loss:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}\bigl{(}\hat{y}_{i}-y_{i}\bigr{)}^{2},$ (1)
where $\hat{y}_{i}$ and $y_{i}$ are the predicted and ground-truth RUL values for the $i$-th sample, respectively. In addition to MSE, we evaluate RMSE and MAE on the validation set to gain further insight into each model’s predictive performance.
Using an Adam optimizer with a batch size of 128 and learning rate $0.001$, we train both the HQRNN and a purely classical RNN for 20 epochs. The two models have an identical number of parameters to ensure a fair comparison.
Each model is trained with 10 different random seeds, and we take the average of the resulting predictions. Surprisingly, as shown in Table 1, the HQRNN outperforms the classical RNN in both RMSE and MAE across all tested parameter configurations. This observation appears to contradict the training-phase results, yet it is consistent with the premise that quantum models may better generalize in data-scarce scenarios caro2022generalization ; schuld2021effect (remember that there are just a $100$ engines in the training dataset).
A possible explanation lies in the quantum circuit’s capacity to capture higher-frequency components of the underlying function xu2019training , which becomes beneficial when the available context (i.e., dataset of engines) is small. In practical terms, our findings indicate that HQRNN may provide robust RUL predictions even when only short segments of sensor readings are accessible. Such resilience to limited context can be especially valuable in real-world aerospace applications, where complete sensor histories are not always available or are expensive to obtain.
Model Mean RMSE Best RMSE Mean MAE Best MAE N of parameters
HQRNN $15.46$ $14.78$ $12.25$ $11.51$ $6793$
RNN-32-16-8-16-32 $16.71$ $15.68$ $13.18$ $\mathbf{12.19}$ $14609$
RNN-20-16-4-8-16 $\mathbf{16.37}$ $15.73$ $\mathbf{12.89}$ $12.51$ $6793$
RNN-16-8-4-8-16 $16.56$ $15.52$ $13.03$ $12.36$ $4233$
RNN-8-4-2-4-8 $29.72$ $\mathbf{15.07}$ $24.52$ $12.20$ $\mathbf{1349}$
Table 1: Comparison of RNN and HQRNN models. RMSE and MAE metrics are computed on the test dataset and averaged over 10 trained models. “Best RMSE” and “Best MAE” columns list the strongest score from any of the 10 runs. Bold values denote the best result among classical models.
V Quantum Circuit Analysis
This section examines the QDI layer used in the QLSTM network through three complementary perspectives: redundancy analysis with ZX calculus, trainability analysis using Fisher Information and expressivity analysis via Fourier analysis.
V.1 Redundancy analysis: ZX Calculus
ZX calculus is a powerful graphical language for representing and simplifying quantum circuits coecke2011interacting ; van2020zx . It employs so-called “spider” nodes, enabling the analysis and optimization of quantum gates through well-defined algebraic rewriting rules. By using ZX techniques, one can potentially reduce the number of parameters and gates in a quantum circuit without changing its overall functionality peham2022equivalence ; wang2024differentiating .
Figure 3 illustrates the original QDI circuit (a) and its optimized form (b). The primary modifications involve rearranging certain weights and merging CNOT gates with them. Despite these adjustments, all $8$ parameters out of $8$ initial ones in the circuit remain essential, indicating that no further simplification is possible without altering the circuit’s behavior. Consequently, this QDI quantum layer structure can be considered ZX-irreducible, implying that it already achieves a high degree of efficiency in terms of parameter usage. Thus none of the weights used here can be considered redundant.
Figure 3: (a) A QDI layer before applying ZX-based parameter reduction. (b) The ZX-reduced QDI layer structure with rearranged weights. Despite these simplifications, no parameters can be removed without affecting the layer’s functionality.
V.2 Trainability Analysis: Fisher Information
In supervised machine learning, one typically trains a model $h_{\theta}(\hat{x})$ on a labeled dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x$ represents input data and $y$ denotes the corresponding labels (targets). The model parameters $\theta=(\theta^{1},\theta^{2},\dots,\theta^{n})$ define a conditional probability distribution $p(y\mid x,\theta)$, which can also be expressed via the joint distribution $p(x,y\mid\theta)$ as
$p(y\mid x,\theta)=\frac{p(x,y\mid\theta)}{p(x)}.$
The set of all such distributions for different $\theta$ forms a manifold =$\mathcal{M}=\{p(y\mid x,\theta),\,\theta\in\Theta\}$.
At each point $\theta$ on this manifold, one can define a tangent space $T_{\theta}\mathcal{M}$, which is locally Euclidean. A common basis for $T_{\theta}\mathcal{M}$ is given by the partial derivatives with respect to each trainable parameter:
$\left\{\frac{\partial}{\partial\theta^{1}},\frac{\partial}{\partial\theta^{2}},\ldots,\frac{\partial}{\partial\theta^{n}}\right\}.$
Using this basis, we define the score function as
$s(\theta)=\nabla_{\theta}\log p(y\mid x,\theta),$
which itself is a tangent vector at $\theta$.
The Fisher Information Matrix (FIM) introduces an inner product on each tangent space, acting as a metric tensor on $\mathcal{M}$:
$g_{\theta}\;:\;T_{\theta}\mathcal{M}\,\times\,T_{\theta}\mathcal{M}\;\longrightarrow\;\mathbb{R}.$
Concretely, the FIM is the covariance matrix of the score amari1998gradient :
$F(\theta)=\mathrm{Cov}\bigl{[}s(\theta)\bigr{]}=\mathbb{E}_{p(x,y\mid\theta)}\bigl{[}s(\theta)\,s(\theta)^{\!\top}\bigr{]}$ (2)
In practice, the FIM is approximated by the sample average.
Moreover, one can evaluate the “volume” of the manifold $\mathcal{M}$ via
[…]
whose logarithm corresponds to the effective dimension introduced in Ref. berezniuk2020scale and serves as a measure of model complexity.
A significant motivation for analyzing the FIM is its connection to the barren plateau phenomenon, wherein gradients vanish for large system sizes mcclean2018barren . A large fraction of near-zero eigenvalues in the FIM often indicates poor trainability, aligning with such plateaus abbas2021power .
Following abbas2021power ; Araz2022Fisher , we compute the FIM of QDI circuit for a Gaussian dataset $x\sim\mathcal{N}(0,1)$. By averaging over all $x$ and $y$, one obtains the mean FIM.
Figure 4 illustrates our main findings. Figure 4 (a) shows the normalized histogram of FIM eigenvalues, highlighting that they are well-distributed with no significant concentration near zero. In Figure 4 (b), the diagonal structure in the averaged FIM indicates that gradients are approximately evenly allocated across parameters, whereas minimal off-diagonal elements suggest weak cross-parameter entanglement and straightforward optimization. Overall, these results confirm that the chosen QDI circuit is well-poised for training in the RUL prediction task.
Figure 4: (a) The normalized histogram of the Fisher eigenvalue spectrum. There are several eigenvalue groups with slightly higher contribution; thus, the circuit doesn’t fully rely on just few significant parameters, especially as all the others follow them closely in terms of eigenvalue percentage. Furthermore, according to abbas2021power , small number of close to zero eigenvalues indicates the resilience to the barren plateau problem. Additionally, none of those groups come close to the $95\%$ majority threshold, thus showing equal contributive distribution. (b) This is an averaged normalized Fisher Information Matrix. The diagonal of this matrix shows that the quantum circuit equally distributes the gradients to all trainable parameters, and there is no evident single-parameter dominance. Anti-diagonal elements aren’t pronounced which means parameters aren’t interconnected resulting in easier optimisation for circuit’s weights.
Figure 5: Real and imaginary parts of the Fourier coefficients for a QDI layer with four input features. The prevalence of non-zero coefficients ($\sim 67\%$) indicates substantial expressivity.
V.3 Expressivity Analysis: Fourier Series
Quantum neural networks employing angle-based encoding can be interpreted through the lens of truncated Fourier series schuld2021effect ; peters2023generalization ; Parfait2024Fourier . Specifically, the circuit’s capacity to represent a function $f(\theta,x)$ can be expressed as a multi-dimensional Fourier series whose degree of truncation depends on the number of encoding repetitions perez2020data . For instance, in a setting with two encoded features, each repeated once, the function becomes:
$f(\theta,x)$ $=\braket{\psi(\theta,x)\mid M\mid\psi(\theta,x)}=$ (3)
$=\sum_{n=-1}^{1}\sum_{m=-1}^{1}c_{nm}(\theta)e^{i(nx_{1}+mx_{2})}$
where $|\psi(\theta,x)\rangle$ is the quantum state of the circuit after all parameterized operations, $M$ denotes the observable, and $c_{nm}$ are complex coefficients determined by the circuit parameters. Although the maximal possible Fourier frequencies are limited by the circuit depth, they can be substantial enough to capture higher-order correlations in real-world datasets. Thus higher number of non-zero coefficients represents more complex dependencies or patterns that model can detect and learn.
In the QDI quantum circuit, we assess Fourier accessibility by encoding four features, each appearing only once in the encoding. Randomly initializing the circuit parameters (using 1000 samples), we then compute the real and imaginary parts of the resulting Fourier spectrum, as shown in Figure 5. Out of $161$ possible frequency components, $109$ are found to have non-negligible amplitude ($\sim 67\%$), signifying that a large portion of the circuit’s Fourier space is indeed accessible. From this perspective, the QDI layer demonstrates considerable expressiveness, which can be advantageous when dealing with complex, high-dimensional data.
V.4 Summary of Circuit Analysis
The ZX-based simplification confirms that the QDI circuit is already optimally parameterized, with no redundant parameters. Meanwhile, the Fisher Information analysis shows that all parameters exhibit strong trainability reinforcing the circuit’s robustness. Finally, the Fourier analysis indicates that a significant portion of the circuit’s Fourier space is accessible, enabling it to represent higher-frequency components crucial for complex time-series data. Taken together, these findings indicate that the circuit’s design is both parameter-efficient and expressively powerful, making it well-suited for the RUL prediction task explored in this study.
VI Discussion
In this study, we presented a HQRNN for predicting RUL of jet engines. Our model employs stacked QLSTM layers, wherein the conventional linear transformations of each LSTM gate are replaced by QDI circuits. As demonstrated in Section V, these QDI layers exhibit strong expressivity, capturing a broad spectrum of frequencies while maintaining a parameter-efficient design.
We tested the HQRNN on the NASA C-MAPSS “FD001” dataset, along with a classical RNN of comparable parameter size. Although the classical model achieved slightly lower losses on the training and validation sets, the HQRNN demonstrated superior generalization on the test set. Specifically, it showed approximately a 5% improvement in both mean RMSE and MAE while also requiring fewer learnable parameters. This suggests that the quantum-enhanced network may be less prone to overfitting and more adept at capturing diverse frequency components that are critical for RUL prediction, especially when data are limited or particularly complex. Analysis of the quantum circuit via ZX calculus, Fisher Information, and Fourier series further confirmed that the chosen QDI circuit is both trainable and expressively powerful, offering a viable path toward improved modeling of real-world degradation patterns.
In comparison with classical ML models (random forest, LASSO regression, etc) and simple ANN models (MLP, CNN, LSTM) on this task, proposed HQRNN model shows the best performance in terms of RMSE error (Table 2).
However, if one wants to achieve state-of-the-art results in such tasks, it is better to use ensembles or combinations of neural network models with addition of complex feature preprocessing (Table 3). For instance, “Auto RUL + LSTM model” asif2022deep uses an improved sophisticated degradation model to assing RUL target labels, while this work sticks to simple piecewise linear degradation model. So it is expected that joint models represented in Table 3 are performing better than our raw HQRNN model. This means that the proposed model can be used not only on its own, but can also be integrated in complex ML pipelines, to further improve their performance.
Type Method RMSE
Classical ML RF zhang2016multiobjective $17.91$
LASSO zhang2016multiobjective $19.74$
SVM zhang2016multiobjective $40.72$
KNR zhang2016multiobjective $20.46$
GB zhang2016multiobjective $15.67$
ANN MLP zhang2016multiobjective $16.78$
CNN sateesh2016deep $18.45$
LSTM zheng2017long $16.14$
Proposed HQRNN $\mathbf{15.46}$
Table 2: Comparison of the proposed HQRNN model with classical ML models and simple neural network models. Best RMSE score is in bold.
Method RMSE
Transformer + TCNN wang2021remaining $12.31$
CNN + LSTM kong2019convolution $16.16$
LSTM + FCLCNN peng2021remaining $11.17$
BLS + TCN yu2021prediction $12.08$
Auto RUL + LSTM asif2022deep $\mathbf{7.78}$
HQRNN (proposed) $15.46$
Table 3: Comparison of the proposed HQRNN model with joint models. Best RMSE score is in bold.
Looking ahead, several promising directions arise from this study. One is to integrate quantum modules into established forecasting algorithms, such as random forests or gradient boosting, where quantum layers can serve as advanced feature encoders. Additionally, transformer-based approaches like TabPFN hollmann2022tabpfn or Chronos ansari2024chronoslearninglanguagetime could benefit from quantum enhancements, potentially improving their ability to model complex temporal dependencies. As quantum hardware evolves, one can also explore adaptive strategies that dynamically vary the circuit size or depth to manage the trade-off between representational capacity and trainability.
By applying a relatively simple quantum-enhanced architecture to a well-known industrial dataset and achieving improvements over classical baselines, this work underscores the practical potential of hybrid quantum-classical models. Although many challenges remain before quantum machine learning reaches widespread industrial use, our results suggest a tangible performance benefit when quantum and classical methods are combined. This outlook could encourage broader adoption of similar hybrid designs and stimulate further research aimed at refining quantum circuit architectures, improving interpretability, and deploying advanced quantum-enhanced solutions for time-series forecasting and other predictive tasks.

### v2

@@
 Each Quantum Long Short-Term Memory gate replaces conventional linear transformations with Quantum Depth-Infused circuits, allowing the network to learn high-frequency components more effectively.
-Experimental results demonstrate that, despite having fewer trainable parameters, the Hybrid Quantum Recurrent Neural Network achieves up to a 5% improvement over a Recurrent Neural Network based on stacked Long Short-Term Memory layers in terms of mean root mean squared error and mean absolute error.
-Moreover, a thorough comparison of our method with established techniques, including Random Forest, Convolutional Neural Network, and Multilayer Perceptron, demonstrates that our approach, which achieves a Root Mean Squared Error of 15.46, surpasses these baselines by approximately 13.68%, 16.21%, and 7.87%, respectively.
-Nevertheless, it remains outperformed by certain advanced joint architectures.
-Our findings highlight the potential of hybrid quantum-classical approaches for robust time-series forecasting under limited data conditions, offering new avenues for enhancing reliability in predictive maintenance tasks.
+Experimental results demonstrate that, despite having fewer trainable parameters, the Hybrid Quantum Recurrent Neural Network achieves up to a $5\%$ improvement over a Recurrent Neural Network based on stacked Long Short-Term Memory layers in terms of mean root-mean-square error and mean absolute error.
+Moreover, a thorough comparison of our method with established techniques, including Random Forest, Convolutional Neural Network, and Multilayer Perceptron, demonstrates that our approach, which achieves a Root Mean Squared Error of $15.46$, surpasses these baselines by approximately $13.68\%$, $16.21\%$, and $7.87\%$, respectively.
+Nevertheless, certain advanced joint architectures still outperform it.
+Our findings highlight the potential of hybrid quantum-classical approaches for robust time-series forecasting under limited-data conditions, offering new avenues for enhancing reliability in predictive maintenance tasks.
 Remaining Useful Life, Quantum Machine Learning, Recurrent Neural Network, LSTM, Predictive Maintenance, Time-Series Forecasting
@@
 I Introduction
-Accurate estimation of the remaining useful life (RUL) of critical machinery is a cornerstone of modern reliability and risk-management strategies si2011remaining ; lee2014prognostics ; berghout2022systematic .
-Nowhere is this more evident than in the aviation industry: the timely prediction of gas turbine jet engine failures not only bolsters safety by preventing catastrophic breakdowns zhang2016multiobjective , but also serves to streamline maintenance and resource allocation kang2021remaining .
+Accurate estimation of the remaining useful life (RUL) of critical machinery is a cornerstone of modern reliability and risk management strategies si2011remaining ; lee2014prognostics ; berghout2022systematic .
+Nowhere is this more evident than in the aviation industry — the timely prediction of gas turbine jet engine failures not only bolsters safety by preventing catastrophic breakdowns zhang2016multiobjective , but also serves to streamline maintenance and resource allocation kang2021remaining .
 Indeed, robust forecasting models enable operators to defer costly repairs until necessary while avoiding the risks of overdue overhauls, ultimately leading to reduced downtime and more efficient fleet utilization huang2019bidirectional ; ferreira2022remaining .
@@
 This development has led to quantum algorithms designed for optimization, simulation, and cryptography, thereby paving the way for more specialized implementations montanaro2016quantum ; Preskill2018quantumcomputingin .
-Finally, the new tool that allows the use of the advantages of ML algorithms and quantum computations emerged — quantum machine learning (QML), which proved its usability by solving problems typified by constrained datasets or high-dimensional feature spaces rebentrost2014quantum ; ciliberto2018quantum ; schuld2018supervised ; cao2017quantum ; Alharbi2024QCNN .
-QML utilises high-dimensional Hilbert spaces to encode input features, allowing it to detect more complex patterns with fewer parameters than its classical counterpart havlivcek2019supervised ; schuld2019quantum .
-This proved to be especially useful when applied to problems with nonstationary or noisy signals, which are frequently encountered in industrial maintenance and prognostic applications  sagingalieva2023photovoltaic ; emmanoulopoulos2022quantummachinelearningfinance ; protasevich2025multi ; lee2025predictive .
-A promising approach for leveraging these quantum advantages involves hybrid quantum-classical neural networks (HQNNs), which combine classical and quantum ML in one single framework arthur2022hybridquantumclassicalneuralnetwork ; sagingalieva2023photovoltaic ; sagingalieva2023hybrid ; Bischof2025Hybrid ; SUN2025130226 .
+Finally, the new tool that allows the use of the advantages of ML algorithms and quantum computations emerged — quantum machine learning (QML), which proved its usability by solving problems typified by constrained datasets or high-dimensional feature spaces rebentrost2014quantum ; ciliberto2018quantum ; schuld2018supervised ; cao2017quantum ; Alharbi2024QCNN ; qml_review_2023 .
+QML utilities high-dimensional Hilbert spaces to encode input features, allowing it to detect more complex patterns with fewer parameters than its classical counterpart havlivcek2019supervised ; schuld2019quantum .
+This proved to be especially useful when applied to problems with nonstationary or noisy signals, which are frequently encountered in industrial maintenance and prognostic applications emmanoulopoulos2022quantummachinelearningfinance ; sedykh2024hybrid ; sagingalieva2025photovoltaic .
+A promising approach for leveraging these quantum advantages involves hybrid quantum-classical neural networks (HQNNs), which combine classical and quantum ML in one single framework kordzanganeh2023parallel ; kordzanganeh2023exponentially ; arthur2022hybridquantumclassicalneuralnetwork ; haboury2024information ; Bischof2025Hybrid ; SUN2025130226 ; patapovich2025superposed .
 Inside this structure, the quantum component plays a specific role in data processing or feature encoding while leaving computationally hard tasks like backpropagation and parameter optimization for the classical routines.
 By selectively introducing quantum effects, HQNNs retain the scalability and reliability of conventional deep learning pipelines broughton2021tensorflowquantumsoftwareframework .
-Preliminary studies indicate that HQNNs can match or exceed the performance of their classical counterparts, often exhibiting greater resilience against overfitting abbas2021power ; Berberich2024training .
-The inherent parameter efficiency of HQNNs further enhances their suitability for scenarios involving complex, high-dimensional inputs or limited data availability — situations frequently observed in predictive maintenance sagingalieva2023hybrid ; anoshin2024hybrid .
+Preliminary studies indicate that HQNNs can match or exceed the performance of their classical counterparts sagingalieva2023hyperparameter ; senokosov2024quantum ; haboury2023supervised ; sagingalieva2025hybrid , often exhibiting greater resilience against overfitting abbas2021power ; Berberich2024training .
+The inherent parameter efficiency of HQNNs further enhances their suitability for scenarios involving complex, high-dimensional inputs or limited data availability — situations frequently observed in predictive maintenance.
 In this paper, we propose a novel Hybrid Quantum Recurrent Neural Network (HQRNN) tailored for jet engine RUL forecasting.
-Our approach replaces the linear transformations in each gate of a classical LSTM with a Quantum Depth-Infused (QDI) circuit sagingalieva2023hybrid ; anoshin2024hybrid ; lusnig2024hybrid , creating a quantum-enhanced LSTM (QLSTM) chen2022quantum ; sagingalieva2023photovoltaic .
+Our approach replaces the linear transformations in each gate of a classical LSTM with a Quantum Depth-Infused (QDI) circuit sagingalieva2023hybrid ; anoshin2024hybrid ; lusnig2024hybrid , creating a quantum-enhanced LSTM (QLSTM) chen2022quantum ; sagingalieva2025photovoltaic .
 By encoding sensor measurements in a quantum Hilbert space, these QDI layers effectively capture complex temporal patterns while requiring fewer trainable parameters than a purely classical baseline chen2022quantum .
 We evaluate our method on the NASA Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset saxena2008turbofan , a widely used benchmark in turbofan engine degradation studies, and demonstrate improved predictive performance in terms of both root mean squared error (RMSE) and mean absolute error (MAE).
-Moreover, experimental results suggest that our hybrid approach can help mitigate the challenges associated with limited data, underscoring its potential benefits in industrial time-series applications sagingalieva2023hybrid .
-The C-MAPSS dataset is a widely used benchmark for predicting RUL of aircraft engines.
+Moreover, experimental results suggest that our hybrid approach can help mitigate the challenges associated with limited data, underscoring its potential benefits in industrial time-series applications sagingalieva2025photovoltaic ; kurkin2025forecasting ; lee2025predictive .
+The C-MAPSS dataset cmapss is a widely used benchmark for predicting RUL of aircraft engines.
 Recent studies addressing this prediction problem include approaches based on Temporal Convolutional Neural Networks (TCNN) combined with Transformers wang2021remaining and joint CNN-LSTM models enhanced with attention mechanisms deng2024prediction ; both articles also provide comprehensive summaries of alternative methods.
@@
 II Dataset
-The NASA C-MAPSS dataset is a widely used multivariate time series for analyzing gas turbine engine degradation.
+The NASA C-MAPSS dataset cmapss is a widely used multivariate time series for analyzing gas turbine engine degradation.
 In this work, we employ the “FD001” subset, which includes run-to-failure data for a fleet of 100 engines.
@@
 In the training set, each engine’s sensor measurements are recorded until failure occurs, whereas in the test set, the measurements end at an arbitrary cycle prior to failure.
-So the goal is to predict the number of remaining operational cycles before failure in the test set, i.e., the number of operational cycles after the last cycle that the engine will continue to operate.
+The goal is to predict the number of remaining operational cycles before failure in the test set, i.e., the number of operational cycles after the last cycle that the engine will continue to operate.
 A vector of true RUL values for the test data is provided.
 We design our predictive model as follows.
-A model gets a fixed size window of consecutive engine cycles as input and yields a single estimated RUL value as output.
+A model gets a fixed-size window of consecutive engine cycles as input and yields a single estimated RUL value as output.
 This means that for each entry in the training dataset there should be a corresponding target RUL value.
 Creation of these RUL values depends on the degradation model.
-A linear degradation model, for instance, linearly decreases the RUL value to zero as engine approaches the point of failure.
-In this work, however, we will use a piecewise linear degradation model, which follows the idea that for some time an engine is operating normally (RUL is constant, also known as “early RUL”) and at some point develops a fault, which leads to linear decrease of its useful life, see Figure 1.
-We chose to set early RUL to 125 cycles, based on the minimum run length of 128 cycles observed in the training data.
+A linear degradation model, for instance, linearly decreases the RUL value to zero as the engine approaches the point of failure.
+In this work, however, we will use a piecewise linear degradation model, which follows the idea that for some time an engine is operating normally (RUL is constant, also known as “early RUL”) and at some point develops a fault, which leads to linear decrease in its useful life, see Figure 1.
+We choose to set early RUL to 125 cycles, based on the minimum run length of 128 cycles observed in the training data.
 Figure 1: Remaining useful life of engine 2 according to linear degradation and piecewise linear degradation models. This work adopts the piecewise linear model with an early RUL threshold of $125$.
@@
 Fig.
-2(b) depicts the internal structure of a single QLSTM layer, highlighting the replacement of each conventional linear gate transformation with a individual QDI layer.
+2(b) depicts the internal structure of a single QLSTM layer, highlighting the replacement of each conventional linear gate transformation with an individual QDI layer.
 Fig.
@@
 Before encoding the input data, a parameterized $R_{x}$ rotation is applied to the initial state of the four qubits using the trained parameters as rotation angles, these are combined with a ring of controlled NOT (CNOT) gates (orange operation block).
-[…]
-$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix}$
+$R_{x}(\theta)=\begin{pmatrix}\cos(\frac{\theta}{2})&-i\sin(\frac{\theta}{2})\\[4.0pt] -i\sin(\frac{\theta}{2})&\ \cos(\frac{\theta}{2})\end{pmatrix}$
+$\mathrm{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix}$
 This is followed by an encoding block represented by $R_{z}$ rotations using input features as rotation angles (blue rectangles).
@@
 In addition to MSE, we evaluate RMSE and MAE on the validation set to gain further insight into each model’s predictive performance.
-Using an Adam optimizer with a batch size of 128 and learning rate $0.001$, we train both the HQRNN and a purely classical RNN for 20 epochs.
+Using an Adam optimizer with a batch size of 128 and a learning rate of $0.001$, we train both the HQRNN and a purely classical RNN for 20 epochs.
 The two models have an identical number of parameters to ensure a fair comparison.
@@
 Surprisingly, as shown in Table 1, the HQRNN outperforms the classical RNN in both RMSE and MAE across all tested parameter configurations.
-This observation appears to contradict the training-phase results, yet it is consistent with the premise that quantum models may better generalize in data-scarce scenarios caro2022generalization ; schuld2021effect (remember that there are just a $100$ engines in the training dataset).
+This observation appears to contradict the training-phase results, yet it is consistent with the premise that quantum models may better generalize in data-scarce scenarios caro2022generalization ; schuld2021effect (remember that there are just $100$ engines in the training dataset).
 A possible explanation lies in the quantum circuit’s capacity to capture higher-frequency components of the underlying function xu2019training , which becomes beneficial when the available context (i.e., dataset of engines) is small.
@@
 Such resilience to limited context can be especially valuable in real-world aerospace applications, where complete sensor histories are not always available or are expensive to obtain.
-Model Mean RMSE Best RMSE Mean MAE Best MAE N of parameters
+Model Mean RMSE Best RMSE Mean MAE Best MAE Number of parameters
 HQRNN $15.46$ $14.78$ $12.25$ $11.51$ $6793$
@@
 V.2 Trainability Analysis: Fisher Information
-In supervised machine learning, one typically trains a model $h_{\theta}(\hat{x})$ on a labeled dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x$ represents input data and $y$ denotes the corresponding labels (targets).
-The model parameters $\theta=(\theta^{1},\theta^{2},\dots,\theta^{n})$ define a conditional probability distribution $p(y\mid x,\theta)$, which can also be expressed via the joint distribution $p(x,y\mid\theta)$ as
+[…]
 $p(y\mid x,\theta)=\frac{p(x,y\mid\theta)}{p(x)}.$
-The set of all such distributions for different $\theta$ forms a manifold =$\mathcal{M}=\{p(y\mid x,\theta),\,\theta\in\Theta\}$.
+The set of all such distributions for different $\theta$ forms a manifold $\mathcal{M}=\{p(y\mid x,\theta),\,\theta\in\Theta\}$.
 At each point $\theta$ on this manifold, one can define a tangent space $T_{\theta}\mathcal{M}$, which is locally Euclidean.
@@
 Moreover, one can evaluate the “volume” of the manifold $\mathcal{M}$ via
-[…]
+$V\,\propto\,\int_{\Theta}\!\sqrt{\det F(\theta)}\,d\theta,$
 whose logarithm corresponds to the effective dimension introduced in Ref. berezniuk2020scale and serves as a measure of model complexity.
@@
 The diagonal of this matrix shows that the quantum circuit equally distributes the gradients to all trainable parameters, and there is no evident single-parameter dominance.
-Anti-diagonal elements aren’t pronounced which means parameters aren’t interconnected resulting in easier optimisation for circuit’s weights.
+Anti-diagonal elements aren’t pronounced which means parameters aren’t interconnected resulting in easier optimization for circuit’s weights.
 Figure 5: Real and imaginary parts of the Fourier coefficients for a QDI layer with four input features. The prevalence of non-zero coefficients ($\sim 67\%$) indicates substantial expressivity.
@@
 VI Discussion
-In this study, we presented a HQRNN for predicting RUL of jet engines.
+In this study, we presented a HQRNN for predicting the RUL of jet engines.
 Our model employs stacked QLSTM layers, wherein the conventional linear transformations of each LSTM gate are replaced by QDI circuits.
@@
 Although the classical model achieved slightly lower losses on the training and validation sets, the HQRNN demonstrated superior generalization on the test set.
-Specifically, it showed approximately a 5% improvement in both mean RMSE and MAE while also requiring fewer learnable parameters.
+Specifically, it showed approximately a $5\%$ improvement in both mean RMSE and MAE while also requiring fewer learnable parameters.
 This suggests that the quantum-enhanced network may be less prone to overfitting and more adept at capturing diverse frequency components that are critical for RUL prediction, especially when data are limited or particularly complex.
 Analysis of the quantum circuit via ZX calculus, Fisher Information, and Fourier series further confirmed that the chosen QDI circuit is both trainable and expressively powerful, offering a viable path toward improved modeling of real-world degradation patterns.
-In comparison with classical ML models (random forest, LASSO regression, etc) and simple ANN models (MLP, CNN, LSTM) on this task, proposed HQRNN model shows the best performance in terms of RMSE error (Table 2).
-However, if one wants to achieve state-of-the-art results in such tasks, it is better to use ensembles or combinations of neural network models with addition of complex feature preprocessing (Table 3).
-For instance, “Auto RUL + LSTM model” asif2022deep uses an improved sophisticated degradation model to assing RUL target labels, while this work sticks to simple piecewise linear degradation model.
-So it is expected that joint models represented in Table 3 are performing better than our raw HQRNN model.
-This means that the proposed model can be used not only on its own, but can also be integrated in complex ML pipelines, to further improve their performance.
+In comparison with classical ML models (random forest, LASSO regression, etc) and simple ANN models (MLP, CNN, LSTM) on this task, the proposed HQRNN model shows the best performance in terms of RMSE error (Table 2).
+However, if one wants to achieve state-of-the-art results in such tasks, it is better to use ensembles or combinations of neural network models with the addition of complex feature preprocessing (Table 3).
+For instance, “Auto RUL + LSTM model” asif2022deep uses an improved sophisticated degradation model to assign RUL target labels, while this work sticks to simple piecewise linear degradation model.
+Therefore, it is expected that joint models represented in Table 3 are performing better than our raw HQRNN model.
+This means that the proposed model can be used not only on its own but also integrated in complex ML pipelines, to further improve their performance.
 Type Method RMSE
@@
 This outlook could encourage broader adoption of similar hybrid designs and stimulate further research aimed at refining quantum circuit architectures, improving interpretability, and deploying advanced quantum-enhanced solutions for time-series forecasting and other predictive tasks.
+Conflict of interest
+The authors declare no conflicts of interest.
+Funding
+No funding was received for this manuscript.